# Mixtures of Gaussian Process Experts with SMC2

### v2

@@ 1 Introduction
Gaussian processes (GP) are a versatile tool for statistical modelling of data where the exact form of the mapping between input and output variables is unknown and have found a plethora of applications. Recent applications of GPs include prediction of battery life times (Richardson et al., 2017), geometry optimization in theoretical chemistry (Denzel and Kästner, 2018), spectral peak structure prediction (Wakabayashi et al., 2018), tokamak plasma modelling (Ho et al., 2019), medical diagnosis using proteomic data (Yu et al., 2022), and pile foundation design (Momeni et al., 2020), just to name a few.
@@ 2 Model
where $(p_{1}(x_{i}\mid\Psi),\dots,p_{K}(x_{i}\mid\Psi))=:{\boldsymbol{p}}(x_{i}\mid\Psi)$ is the gating network which takes values in the $K-1$-dimensional simplex and the expert $p(y\mid x_{i},\Theta_{k})$ is a conditional model, with local relevance determined by $p_{k}(x_{i}\mid\Psi)$ at the input location $x_{i}$. Here, $\Psi:=(\psi_{1},\ldots,\psi_{K})$ is a shorthand for the gating network parameters and $\Theta=(\Theta_{1},\ldots,\Theta_{K})$ contains the expert parameters. MoEs can be augmented with a set of latent allocation variables $c_{i}\in\{1,\ldots,K\}$ and equivalently represented by
@@ 2.1 Experts
[…]
@@ 2.2 Gating networks
To ensure the gating network $\boldsymbol{p}(x_{i})=(p_{1}(x_{i}),\dots,p_{K}(x_{i}))^{T}$ takes values in the $K-1$-dimensional simplex, we assume $0<\mathcal{K}_{k}(x\mid\psi)<\infty$ for all $x\in\mathbb{R}^{D}$. Examples of kernels used in literature include the multinomial logistic model with $\mathcal{K}_{k}(x\mid\psi_{k})=\exp(\psi_{k}^{T}x)$ (Li et al., 2011) and the factorized kernel $\mathcal{K}_{k}(x\mid\psi_{k})=v_{k}\prod_{d=1}^{D}\mathcal{K}_{k,d}(x_{d}\mid\psi_{k,d})$ (Antoniano-Villalobos et al., 2014). In the latter, mixed types of inputs are allowed by specifying each $\mathcal{K}_{k,d}(x_{d}\mid\psi_{k,d})$ to correspond to the kernel of standard distributions, depending on the nature of the $d$th input. We emphasize that this does not correspond to any model assumptions on the input, which in fact may be fixed.
@@ 4 Inner SMC sampler
$p_{t}(Y\mid X,C):=\int p(Y\mid X,C,\Theta)^{\kappa^{(t)}}\pi_{0}(\Theta)\,\text{d}\Theta=\prod\limits_{k=1}^{K}\int p\left(Y_{k}\mid X_{k},\Theta_{k}\right)^{\kappa^{(t)}}\pi_{0}(\Theta_{k})\,\text{d}\Theta_{k},$ (22)
and where $t$ stands for the “time” step of the SMC or SMC2 sampler and $\kappa^{(t)}$ is a strictly increasing sequence of parameters, $10=\kappa(0)<\dots<\kappa(t-1)<\kappa(t)<\dots<\kappa(T)=$ 1, controlling the degree of tempering. We discuss the explicit construction of the tempering sequence, which depends on the outer SMC sampler, in the following Section.
@@
is a non-negative and unbiased estimator (Moral, 2004; Dai et al., 2022) of $p_{t}(Y\mid X,C)$ in Eq. (22) with $\Theta^{(s)}_{m}$ being the output from an SMC with fixed $C$. In other words, the initial particles $\Theta^{(0)}_{m}$ are i.i.d from the prior distribution $\Theta^{(0)}_{m}\sim\pi_{0}(\Theta)$. For subsequent iterations $t=1,\dots,T$, the particles are obtained as $\Theta^{(t)}_{m}\sim\mathcal{M}_{t}(\widehat{\,\Theta}^{(t)}_{m},\,\cdot)$, where $\mathcal{M}_{t}$ is an MCMC kernel, $\widehat{\Theta}^{(t)}_{m}=\Theta_{I_{m}^{(t)}}^{(t)}$ are resampled particles, and $I_{m}^{(t)}\sim\text{Categorical}({\bf\omega}^{(t)})$ are resampling indices for $m=1,\dots,M$, with ${\bf\omega}^{(t)}=(\omega^{(t)}_{1},\dots,\omega^{(t)}_{M})$ and the individual particle weight are given as $\omega^{(t)}_{m}\propto p(Y\mid X,C,\Theta^{(t-1)}_{m})^{\kappa^{(t)}-\kappa^{(t-1)}}$. Here $\mathcal{M}_{t}$ is an MCMC kernel which keeps
@@
Estimation of $\{(\Theta,\omega)_{m}\}_{m=1}^{M}\approx p_{t}(\Theta\mid X,Y)\propto p(Y\mid X,\Theta)^{\kappa(t)}\pi_{0}(\Theta)$ and $\pi(Y\mid X)$:
while $\kappa(s)<\kappa(t)$ do
@@
     Define $\omega^{(s)}_{m}:=p(Y\mid X,\Theta_{m}^{(s-1)})^{\kappa(s)-\kappa(s-1)}$
     Select new indices $I_{m}^{(s)}$ according to $\omega^{(s)}_{m}$
@@
     Set $\widehat{Z}_{s}^{M}\leftarrow\left(\frac{1}{M}\sum_{m=1}^{M}\omega^{(s)}_{m}\right)\widehat{Z}_{s-1}^{M}$, following Eq. (25).
     Draw $\Theta^{(s)}_{m}\sim\mathcal{M}_{s}(\Theta^{(s)}_{I_{m}^{(s)}},\,\cdot)$, where $\mathcal{M}_{s}$ is an MCMC kernel targeting $p_{s}(\Theta\mid X,Y)$.
@@ 6 Predictive distribution
     Resample $U^{(t)}_{j}$ given particles weights $\boldsymbol{w}^{(t)}$, where $w_{j}^{(t)}$ follow Eq. (30).
@@
     Draw $U^{(t)}_{j}\sim\mathcal{K}_{t}(\widehat{U}^{(t)}_{j},\,\cdot)$ – PMCMC kernel (Algorithm 2) targeting $\pi_{t}$ in Eq. (27).
@@ 7 Numerical examples
Figure 9: First synthetic data set: non-stationary function, homoskedastic noise. In a), the generated data points are shown (in blue) with a 90% highest density region of the underlying ground truth distribution in black and the true function as a solid red line. In b), c), and d) the highest density region for MoE with SMC2 is shown for different values of the Dirichlet concentration parameter $\alpha=0.1,1,K/2$. In e), f), and g) the highest density regions are shown for IS-MoE with the same Dirichlet concentration parameters. In the bottom row, h), i), and j) show the predictive distributions for a single GP, RBCM, and treedGP, respectively.
Figure 10: Second synthetic data set: discontinuous function, heteroskedastic noise. In a) the generated data points are shown (in blue) with a highest density region of the underlying ground truth distribution in black and the true function as a solid red line. In b), c), and d) the highest density region for MoE with SMC2 is shown for different values of the Dirichlet concentration parameter $\alpha=0.1,1,K/2$. In e), f), and g) the highest density regions are shown for IS-MoE with same Dirichlet concentration parameters. In the bottom row, h), i), and j) show the highest density regions given for a single GP, RBCM, and treedGP, respectively.
$L^{1}$ $L^{2}$ $L^{1}$ median LL
@@
Table 2: First synthetic data set: non-stationary function, homoskedastic noise. Column $L^{1}$ shows the $L^{1}$ vector norm of the difference between the estimated distributions and the ground truth distribution used to generate the data set for a given method. Column $L^{2}$ shows the $L^{2}$ vector norm of the difference between the aforementioned distributions and column $L^{1}$ median shows the $L^{1}$ vector norm of the distance between the posterior predictive median and the ground truth median. Column LL displays predictive log-likelihood results.
$L^{1}$ $L^{2}$ $L^{1}$ median LL
@@
Table 3: Second synthetic data set: discontinuous function, heteroskedastic noise. Column $L^{1}$ shows the $L^{1}$ vector norm of the difference between the estimated distributions and the ground truth distribution used to generate the data set for a given method. Column $L^{2}$ shows the $L^{2}$ vector norm of the difference between the aforementioned distributions and column $L^{1}$ median shows the $L^{1}$ vector norm of the distance between the posterior predictive median and the ground truth median. Column LL displays predictive log-likelihood results.
Figure 11: First synthetic data set: non-stationary function, homoskedastic noise. On the left, the highest density regions and median for SMC2-MoE are shown in black and red, along with the data points in blue, for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=0.1$, $\alpha=1$, and $\alpha=K/2$ are used and correspond to the top, middle, and bottom rows, respectively.
Figure 12: Second synthetic data set: discontinuous function, heteroskedastic noise. On the left, the highest density regions and median for SMC2-MoE are shown in black and red, along with the data points in blue, for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=0.1$, $\alpha=1$, and $\alpha=K/2$ are used and correspond to the top, middle, and bottom rows, respectively.
@@
Figure 13: Motorcycle data. On the left, the highest density regions and median are shown in black and red along with the data points in blue for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=0.1$, $\alpha=1$, and $\alpha=K/2$ are used and correspond to the top, middle, and bottom plots, respectively.
Figure 14: 3D NASA Langley glide-back booster simulation data. On the left, the highest density regions and median for a one-dimensional slice of the input space are shown in black and red along with the data points in blue for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=2.7$, $\alpha=6.75$, and $\alpha=K/2=13.5$ are used and correspond to the top, middle, and bottom plots, respectively.
@@
Figure 15: 4D Colorado precipitation data. On the left, the highest density regions and median for a one-dimensional slice of the input space are shown in black and red along with the data points in blue for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=1.6$, $\alpha=4$, and $\alpha=K/2=8$ are used and correspond to the top, middle, and bottom plots, respectively.

### v3

@@ 1 Introduction
Gaussian processes (GP) are a versatile tool for statistical modelling of data where the exact form of the mapping between input and output variables is unknown and have found a plethora of applications. Recent applications of GPs include prediction of battery life times (Richardson et al., 2017), geometry optimization in theoretical chemistry (Denzel and Kästner, 2018), spectral peak structure prediction (Wakabayashi et al., 2018), tokamak plasma modelling (Ho et al., 2019), medical diagnosis using proteomic data (Yu et al., 2022), pile foundation design (Momeni et al., 2020), and computational biology (Otto et al., 2024), just to name a few.
@@ 2 Model
where $(p_{1}(x_{i}\mid\Psi),\dots,p_{K}(x_{i}\mid\Psi))=:{\boldsymbol{p}}(x_{i}\mid\Psi)$ is the gating network which takes values in the K1K-1-dimensional simplex and the expert $p(y\mid x_{i},\Theta_{k})$ is a conditional model, with local relevance determined by $p_{k}(x_{i}\mid\Psi)$ at the input location $x_{i}$. Here, $\Psi:=(\psi_{1},\ldots,\psi_{K})$ is a shorthand for the gating network parameters and $\Theta=(\Theta_{1},\ldots,\Theta_{K})$ contains the expert parameters. MoEs can be augmented with a set of latent allocation variables $c_{i}\in\{1,\ldots,K\}$ and equivalently represented by
@@ 2.1 Experts
where $x_{i},x_{j}\in X_{k}$, $\sigma_{k,f}^{2}$ is the signal variance, and $l_{d}$ is the length scale along the $d$th dimension, with $\delta(x)$ denoting the Kronecker delta function and $\theta_{k}=(\sigma_{k,\varepsilon},\sigma_{k,f},l_{k,1},\ldots,l_{k,D})$. Furthermore, we use the notation $\Theta_{k}=(m_{k},\theta_{k})$ for the kthth GP parameters and $\Theta=(\Theta_{1},\ldots,\Theta_{K})$ for all the GP parameters. An example partition and its corresponding GP fits are illustrated in Figure 1.
@@ 2.2 Gating networks
To ensure the gating network $\boldsymbol{p}(x_{i})=(p_{1}(x_{i}),\dots,p_{K}(x_{i}))^{T}$ takes values in the K1K-1-dimensional simplex, we assume $0<\mathcal{K}_{k}(x\mid\psi)<\infty$ for all $x\in\mathbb{R}^{D}$. Examples of kernels used in literature include the multinomial logistic model with $\mathcal{K}_{k}(x\mid\psi_{k})=\exp(\psi_{k}^{T}x)$ (Li et al., 2011) and the factorized kernel $\mathcal{K}_{k}(x\mid\psi_{k})=v_{k}\prod_{d=1}^{D}\mathcal{K}_{k,d}(x_{d}\mid\psi_{k,d})$ (Antoniano-Villalobos et al., 2014). In the latter, mixed types of inputs are allowed by specifying each $\mathcal{K}_{k,d}(x_{d}\mid\psi_{k,d})$ to correspond to the kernel of standard distributions, depending on the nature of the $d$th input. We emphasize that this does not correspond to any model assumptions on the input, which in fact may be fixed.
@@ 4 Inner SMC sampler
$p_{t}(Y\mid X,C):=\int p(Y\mid X,C,\Theta)^{\kappa^{(t)}}\pi_{0}(\Theta)\,\text{d}\Theta=\prod\limits_{k=1}^{K}\int p\left(Y_{k}\mid X_{k},\Theta_{k}\right)^{\kappa^{(t)}}\hskip-2.84544pt\pi_{0}(\Theta_{k})\,\text{d}\Theta_{k},$ (22)
and where $t$ stands for the “time” step of the SMC or SMC2 sampler and $\kappa^{(t)}$ is a strictly increasing sequence of parameters, $0=\kappa^{(0)}<\dots<\kappa^{(t-1)}<\kappa^{(t)}<\dots<\kappa^{(T)}=1$, controlling the degree of tempering. We discuss the explicit construction of the tempering sequence, which depends on the outer SMC sampler, in the following Section.
@@
[…]
@@
Estimation of $\{(\Theta,\omega)_{m}\}_{m=1}^{M}\approx p_{t}(\Theta\mid X,Y)\propto p(Y\mid X,\Theta)^{\kappa^{(t)}}\pi_{0}(\Theta)$ and $\pi(Y\mid X)$:
while $\kappa^{(s)}<\kappa^{(t)}$ do
@@
  Define $\overline{\omega}^{(s)}_{m}:=p(Y\mid X,\Theta_{m}^{(s-1)})^{\kappa^{(s)}-\kappa^{(s-1)}}$
  Set $\widehat{Z}_{s}^{M}\leftarrow\left(\frac{1}{M}\sum_{m=1}^{M}\overline{\omega}^{(s)}_{m}\right)\widehat{Z}_{s-1}^{M}$, following Eq. (25).
  Select new indices $I_{m}^{(s)}$ according to $\omega^{(s)}_{m}\propto\overline{\omega}^{(s)}_{m}$
  Draw $\Theta^{(s)}_{m}\sim\mathcal{M}_{s}(\Theta^{(s-1)}_{I_{m}^{(s)}},\,\cdot)$, where $\mathcal{M}_{s}$ is an MCMC kernel targeting $p_{s}(\Theta\mid X,Y)$.
@@ 6 Predictive distribution
  Resample $U^{(t-1)}_{j}$ given particles weights $\boldsymbol{w}^{(t)}$, where $w_{j}^{(t)}$ follow Eq. (30).
@@
[…]
@@ 7 Numerical examples
Figure 9: First synthetic data set: non-stationary function, homoskedastic noise. In a), the generated data points are shown (in blue) with a 90% highest density region of the underlying ground truth distribution in gray and the true function as a solid black line. In b), c), and d) the highest density region and posterior median for MoE with SMC2 are shown for different values of the Dirichlet concentration parameter $\alpha=0.1,1,K/2$. In e), f), and g) similar results are shown for IS-MoE with the same Dirichlet concentration parameters. In the bottom row, h), i), and j) show the predictive distributions for a single GP, RBCM, and treedGP, respectively.
Figure 10: Second synthetic data set: discontinuous function, heteroskedastic noise. In a) the generated data points are shown (in blue) with a highest density region of the underlying ground truth distribution in gray and the true function as a solid black line. In b), c), and d) the highest density region and posterior median for MoE with SMC2 are shown for different values of the Dirichlet concentration parameter $\alpha=0.1,1,K/2$. In e), f), and g) similar results are shown for IS-MoE with same Dirichlet concentration parameters. In the bottom row, h), i), and j) show the predictive distributions for a single GP, RBCM, and treedGP, respectively.
$L^{1}$ $L^{2}$ $L^{1}$ median $L$
@@
Table 2: First synthetic data set: non-stationary function, homoskedastic noise. Column $L^{1}$ shows the $L^{1}$ vector norm of the difference between the estimated distributions and the ground truth distribution used to generate the data set for a given method. Column $L^{2}$ shows the $L^{2}$ vector norm of the difference between the aforementioned distributions and column $L^{1}$ median shows the $L^{1}$ vector norm of the distance between the posterior predictive median and the ground truth median. Column $L$ displays predictive log-likelihood results.
$L^{1}$ $L^{2}$ $L^{1}$ median $L$
@@
Table 3: Second synthetic data set: discontinuous function, heteroskedastic noise. Column $L^{1}$ shows the $L^{1}$ vector norm of the difference between the estimated distributions and the ground truth distribution used to generate the data set for a given method. Column $L^{2}$ shows the $L^{2}$ vector norm of the difference between the aforementioned distributions and column $L^{1}$ median shows the $L^{1}$ vector norm of the distance between the posterior predictive median and the ground truth median. Column $L$ displays predictive log-likelihood results.
Figure 11: First synthetic data set: non-stationary function, homoskedastic noise. On the left, the highest density regions and median for SMC2-MoE are shown in gray and black, along with the data points in blue, for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=0.1$, $\alpha=1$, and $\alpha=K/2$ are used and correspond to the top, middle, and bottom rows, respectively.
Figure 12: Second synthetic data set: discontinuous function, heteroskedastic noise. On the left, the highest density regions and median for SMC2-MoE are shown in gray and black, along with the data points in blue, for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=0.1$, $\alpha=1$, and $\alpha=K/2$ are used and correspond to the top, middle, and bottom rows, respectively.
@@
Figure 13: Motorcycle data. On the left, the highest density regions and median are shown in gray and black along with the data points in blue for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=0.1$, $\alpha=1$, and $\alpha=K/2$ are used and correspond to the top, middle, and bottom plots, respectively.
Figure 14: 3D NASA Langley glide-back booster simulation data. On the left, the highest density regions and median for a one-dimensional slice of the input space are shown in gray and black along with the data points in blue for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=2.7$, $\alpha=6.75$, and $\alpha=K/2=13.5$ are used and correspond to the top, middle, and bottom plots, respectively.
@@
Figure 15: 4D Colorado precipitation data. On the left, the highest density regions and median for a one-dimensional slice of the input space are shown in gray and black along with the data points in blue for different values of the Dirichlet concentration parameter $\alpha$. In the middle column, the posterior distribution for the number of non-empty experts is shown. The last column shows the posterior similarity matrices summarizing the posterior over partitions. Values of $\alpha=1.6$, $\alpha=4$, and $\alpha=K/2=8$ are used and correspond to the top, middle, and bottom plots, respectively.